OpenStudy (anonymous):

Solve. 7^x - 5 = 7^4x + 7 x = -4 x =-12/5 x = 2/3 x = 4

OpenStudy (anonymous):

on the right side, is X exponent too or only 4 ?

OpenStudy (anonymous):

4x+7 ..all of it is an exponent .

OpenStudy (anonymous):

how about the left side ?

OpenStudy (anonymous):

x-5 is the exponent

OpenStudy (anonymous):

\[\LARGE 7^{x-5}=7^{4x+7} \quad \quad \Longrightarrow x-5=4x+7\] can you do it now ? :)

OpenStudy (anonymous):

x-5=4x+7 start from here..